OpenStudy (anonymous):

Write the equation y^2+4y+4x=3 in standard form and graph. Name the curve, find the center, focus and directrix. Help me solve please!

OpenStudy (anonymous):

bring equation in form of (y-h)^2=(x-k)+c where h,k is centre

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

y^2+4y=-4x+3 add 2on both sides y^2+4y+2=-4x+3+2 (y+2)^2=-4x+5 (y-(-2))^2=-4x+4+1=-4(x-1)+1 centre is -2,-1

OpenStudy (anonymous):

why did u add 2

OpenStudy (anonymous):

to bring y^2+4y+2 into perfect square form because (a+b)^2=a2+2ab+b2 =y^2+2*2*y+2